### 24 Jan 2024 v1

# APT-Pipe: An Automatic Prompt-Tuning Tool for Social Computing Data Annotation

Abstract.
Recent research has highlighted the potential of LLM applications, like ChatGPT, for performing label annotation on social computing text. However, it is already well known that performance hinges on the quality of the input prompts. To address this, there has been a flurry of research into prompt tuning — techniques and guidelines that attempt to improve the quality of prompts. Yet these largely rely on manual effort and prior knowledge of the dataset being annotated. To address this limitation, we propose APT-Pipe, an automated prompt-tuning pipeline. APT-Pipe aims to automatically tune prompts to enhance ChatGPT’s text classification performance on any given dataset. We implement APT-Pipe and test it across twelve distinct text classification datasets. We find that prompts tuned by APT-Pipe help ChatGPT achieve higher weighted F1-score on nine out of twelve experimented datasets, with an improvement of 7.01% on average. We further highlight APT-Pipe’s flexibility as a framework by showing how it can be extended to support additional tuning mechanisms.
Human computation, crowdsourcing, prompt-tuning, large language models, data annotation
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; May 13–17, 2024; Singapore
1. Introduction
Data annotation commonly relies on human intelligence and suffers from challenges related to annotator quality, data comprehension, domain knowledge, and time cost (Haq et al., 2022; Huang et al., 2016; Veit et al., 2016; Chalkidis et al., 2020). Recent research has highlighted that prompt-based LLM applications, like ChatGPT, have the potential to substitute human intelligence as a data annotator for social computing tasks, e.g., sentiment analysis (Bang et al., 2023; Zhu et al., 2023) and hate speech detection (Huang et al., 2023). Use of tools like ChatGPT for such annotation tasks can significantly improve the cost and speed for researchers. However, such models are heavily impacted by prompt quality, whereby poorly tuned prompt can severely damage annotation accuracy (Shum et al., 2023). This creates the need for tools that can help researchers design and evaluate effective prompts.
To date, most prompt-tuning for annotation tasks is done either manually (Liao et al., 2022) or by injecting prior knowledge of the dataset (Gilardi et al., 2023; Huang et al., 2023). Prompt-tuning for ChatGPT, however, usually introduces demands on expertise that can be hard to address. Thus, we argue that an automated prompt-tuning tool for ChatGPT tasks could remove a significant manual load related to learning prior knowledge of the dataset to annotate. Researchers have already begun to propose techniques for such automation (Petroni et al., 2019; Min et al., 2021; Bang et al., 2023). We focus on three early-stage ideas: introducing standardized templates that (almost) guarantee consistent responses; introducing few-shot examples into prompts to guide annotations; and embedding additional text metadata to augment raw text information. In practice, however, implementing these simple concepts raises key challenges. First, ChatGPT can generate unpredictable and unstructured responses that make parsing annotations difficult (as its responses lack a standard format). Although cloze-style prompts can improve responses’ parsability, such templates usually introduce a trade-off on ChatGPT’s time efficiency (Petroni et al., 2019). Second, although preliminary evidence suggests that performance can be improved by including few-shot examples of annotation (Min et al., 2021), manual exemplar selection can be time consuming. Thus, techniques need to be developed to automatically identify suitable few-shot examples for the prompt. Third, although embedding natural language processing (NLP) metrics in prompts (Bang et al., 2023) has been shown to increase performance, the selection of such metrics requires domain specific expertise. Thus, techniques need to be developed to select the best permutations of NLP metrics to include in prompts. Importantly, the suitability of the above approaches may differ on a per-dataset basis, making a single standardized approach impractical.
To address these challenges, we propose the Automatic Prompt-Tuning Pipeline (APT-Pipe), an extensible prompt-tuning framework for improving text-based annotation task performance. APT-Pipe only requires humans to annotate a small sample subset of data to offer a robust ground truth. It then exploits this annotated subset to automatically tune and test prompts with the goal of improving ChatGPT’s classification performance. APT-Pipe addresses the above three challenges using a three-step prompt-tuning pipeline. It first generates JSON template prompts that can make ChatGPT respond consistently in a desired format. APT-Pipe then automatically computes the most suitable few-shot exemplars to include in prompts. Finally, through a number of iterations, APT-Pipe tests a variety of prompt configurations with NLP metrics to identify which work best for the particular dataset. These three steps are derived from our experience and state-of-the-literature. However, we acknowledge that many other prompt-tuning techniques are likely to emerge in the coming years. Consequently, we design APT-Pipe in an extensible and modular fashion, allowing other researchers to easily “plug-in” additional tuning strategies.
We implement APT-Pipe with the ChatGPT gpt-3.5-turbo model and evaluate it across twelve distinct text classification datasets drawn from three domains in social computing. APT-Pipe will be made publicly available for use. Our results shows that prompts tuned by APT-Pipe can improve ChatGPT’s classification performance on nine out of twelve datasets, with a 7.01% increase of weighted F1-score on average. Moreover, prompts tuned by APT-Pipe can decrease ChatGPT’s computing time for classification, while providing over 97% of all responses in a consistent and easy-to-parse format. Finally, we highlight APT-Pipe’s flexibility by extending its framework with two state-of-art tuning mechanisms called Chain-of-Thought and Tree-of-Thought.
The contributions of this paper can be summarized as follows:
  • We propose APT-Pipe, an automatic and extensible prompt-tuning pipeline to improve ChatGPT’s performance for social computing data annotation.
  • We implement and validate APT-Pipe’s effectiveness. We show that APT-Pipe can improve ChatGPT’s classification performance on nine out of twelve different datasets, with improvement on ChatGPT’s F1-score by 7.01% on average.
  • We show that APT-Pipe prompts generate annotations in a consistent format and take less time compared to existing prompt designs. On average, 97.08% responses generated by APT-Pipe are parsable by desired format and it decreases more than 25% time cost of the baselines.
  • We highlight APT-Pipe’s extensibility by introducing two additional prompt-tuning techniques, Chain of Thought (CoT) and Tree of Thought (ToT). We show that CoT improves APT-Pipe’s F1-scores on five datasets by 4.01%, and ToT improves APT-Pipe’s F1-scores on six datasets by 3.49%. Importantly, we show that APT-Pipe can learn these improvements on a per-dataset basis, automatically tuning the prompts without human intervention.
2. Background & Related Work
In this section, we present a review of the literature on LLMs’ applications in social computing and prompt-tuning for text annotation.
2.1. LLMs in Social Computing
Social computing research relies extensively on human-annotated datasets. Such annotations are time consuming and often costly. The release of LLM tools such as ChatGPT (Bang et al., 2023; Guo et al., 2023) has uncovered a range of possibilities to automate text data labeling tasks due to their robust performance in natural language comprehension and reasoning (Kojima et al., 2022).
Recently, some studies (Sallam et al., 2023; Huang et al., 2023) have employed LLMs to assist in text annotation in social computing. Sallam et al. (Sallam et al., 2023) employ LLMs to discriminate misinformation about vaccines on social media platforms. Huang et al. (Huang et al., 2023) exploit LLMs to annotate implicit hateful tweets and report nearly 80% accuracy. Zhu et al. (Zhu et al., 2023) evaluate the potential of LLMs to annotate five different types of text data. However, work has revealed that the annotation performance of LLMs varies across different LLMs and different domains and datasets (Aiyappa et al., 2023; Shen et al., 2023). The above literature demonstrates that, although it is feasible to employ LLMs for text annotation tasks, the generality and reliability of LLMs need to be improved.
2.2. Prompt-Tuning for Text Annotation
In existing LLM-based text annotation methods (Alizadeh et al., 2023; Luo et al., 2023; Kuzman et al., 2023; Reiss, 2023), several prompt-tuning techniques have been utilized to enhance the performance of LLMs. Alizadeh et al. (Alizadeh et al., 2023) investigate the performance difference between LLMs and crowd-workers in text annotation tasks. They employ natural language formatted prompts in zero-shot tests. For few-shot tests, they employ the Chain-of-Thought (CoT) (Wei et al., 2022) prompting technique, where LLMs are provided with the question and step-by-step reasoning answers as examples. Kuzman et al. (Kuzman et al., 2023) conduct genre identification tasks using LLMs by feeding prompts with manually defined task descriptions, and manually extracting the answers from responses. Reiss (Reiss, 2023) applies LLMs to text annotation tasks by manually designing 10 prompts, feeding them to LLMs, and evaluating their performance. From the above literature, we observe that existing prompt-tuning methods require the participation of domain expertise and human judgement, which hinders the utilization of LLMs in automated annotation tasks. Using ChatGPT as a case-study, we focus on proposing a pipeline that can automatically perform prompt tuning.
3. APT-Pipe Design
We now describe our pipeline for tuning a text classification prompt. The pipeline’s goal is to generate a tuned prompt to improve ChatGPT’s classification performance on a per-dataset basis. The pipeline operates in discrete steps, and we have designed our framework in an extensible fashion, allowing future modifications to be “plugged-in” as new steps. The current implementation has three steps. First, the prompt engineer must personalize a JSON-based prompt following our template (Step 1). Following this, the pipeline searches for suitable few-shot learning examples to include in the prompt (Step 2) and augmentary text metrics that can guide ChatGPT’s reasoning (Step 3). To inform this decision, APT-Pipe iteratively tests prompt variations on a small subset of manually annotated data (typically a few hundred samples) to learn the best combination of information to include in the final prompt. Once complete, the prompt can be used on the remaining dataset.
3.1. Step 1: Preparing Initial Prompt
Figure 1. Flow chart of Step 1. Prompt on the right-hand side shows example text for clickbait news headline detection.
Figure 1 illustrates the process to prepare the initial prompts. The prompt engineer must first input an initial prompt using a JSON syntax. We use JSON because encoded notations can facilitate language models’ ability to understand and respond with more precise answers (Zhou et al., 2022; Yu et al., 2023a). Moreover, designing prompts as JSONs is convenient for prompt augmentation, as APT-Pipe can easily embed additional information using suitable key-value pairs (Yu et al., 2023a; Kumar et al., 2023). Note, we have also experimented with various commonly applied prompt templates, i.e. natural language (Bang et al., 2023), dictionary (Zhu et al., 2023), cloze question (Hu et al., 2021), and JSON (Zhuo et al., 2023). Based on our findings, we observe that JSON prompts generate more precise responses in a consistent format. Although the engineer could input any preferred JSON template, by default, we rely on the following template:
xxxx‘‘Prompt’’: ‘‘Classify the following text by given labels for specified xxxxtask.’’,
xxxx‘‘Text’’: ‘‘{text to classify}’’,
xxxx‘‘Task’’: ‘‘{task domain}’’,
xxxx‘‘Labels’’: [‘‘Label 1’’, ‘‘Label 2’’, ...],
xxxx‘‘Desired format’’: {
xxxxxxxx‘‘Label’’: ‘‘<label_for_classification>’’
xxxx}
This prompt template must then be populated with values by the user. The initial sentence clarifies that the prompt pertains to text classification, with “Text” serving as a placeholder for the input text that the user wishes to annotate. Naturally, each individual text item will be annotated one-by-one using a separate prompt invocation. “Task” specifies the domain for classification (e.g., stance detection), and “Labels” enumerates the potential labels for classifying the text. The user must fill these out to reflect the specific annotation task specification. Finally, the “Desired Format” specifies the format that ChatGPT should utilize to form its answers. Afterwards, the generated classification prompts in Step 1 are treated as inputs for the pipeline to conduct prompt-tuning in Step 2 next.
3.2. Step 2: Prompt-tuning with Few-shot
Figure 2. Flow chart of Step 2.
In this step, the pipeline tunes the prompt taken from Step 1 by employing a few-shot learning approach. This setting gives the model a few demonstrations of the correct annotations at inference time. Such a technique enables the model to study ”in-context” knowledge from given examples and provide better decisions (Brown et al., 2020). We embed this into the existing JSON prompt text by including a set of samples from the dataset alongside their ground-truth labels. The challenge here is identifying the best samples to include in the few-shot examples. Figure 2 presents an overview and we detail the process as follows.
APT-Pipe first conducts exemplar selection. To select the samples to include from the manually annotated set, we rely on the method proposed by Yang et al. (Yang et al., 2022). Specifically, we select the text samples from the dataset with the highest textual similarity with the text to classify. Thus, we use OpenAI’s text-embedding-ada-002 (Neelakantan et al., 2022) model to generate the embeddings for each data sample in the dataset and then calculate the pairwise cosine similarity between the embedding of each exemplar and the embedding of the text to classify. Consequently, we focus on the top $n$ exemplars with the highest cosine similarity. Note, $n$ is configurable, allowing the user to select a range of potential values to experiment on. By default, $n=5$ as a five-shot setting is commonly used in related works (Han et al., 2022; Ma et al., 2021). We include the $n$ examples in the template as follows:
‘‘Examples’’: [
xxxx{‘‘Text’’: ‘‘{exemplar 1}’’, ‘‘Label’’: ‘‘{exemplar 1’s ground-truth xxxxlabel}’’},
xxxx{‘‘Text’’: ‘‘{exemplar 2}’’, ‘‘Label’’: ‘‘{exemplar 2’s ground-truth xxxxlabel}’’},...]
APT-Pipe then uses the small data subset, manually annotated by the prompt engineer, to test if the few shot additions improve performance. First, APT-Pipe requests ChatGPT to annotate the subset without any few-shot examples included; this serves as the baseline. APT-Pipe then repeats the process, including the $n$ few-shot examples (note, $n$ can cover a range of values). For all configurations, APT-Pipe calculates the weighted F1-score of the classification, by comparing ChatGPT’s outputs to the human ground truth labels. We do this check because prior work has shown that including few-shot examples can actually degrade performance on certain datasets (Perez et al., 2021; Zhao et al., 2021). If the F1-score does not drop, APT-Pipe outputs the prompt that attained the highest F1 score improvements. If the F1-score drops, we remove the key-value pairs of the examples and then pass the original prompt to Step 3.
3.3. Step 3: Prompt-tuning with NLP Metrics
Figure 3. Flow chart of Step 3.
As highlighted in (Hu et al., 2021; Ye et al., 2022), domain-relevant text knowledge can play a helpful role in prompt-tuning. Therefore, in Step 3, APT-Pipe tunes the prompt by augmenting it with additional metadata that describes the text to annotate. Figure 3 presents an overview, as well as a pseudo-code representation in Algorithm 1. For this, APT-Pipe first calculates a pre-defined set of NLP metrics for all items of the dataset. Our implementation currently includes sentiment, emotion, toxicity, and topic metrics. We emphasize that our pipeline is extensible and can include any other metadata.
To prepare for Step 3, we initially divide the data annotated by the engineer into two segments: a training set comprising 75% of the data, and a validation set consisting of the remaining 25%. Following this, we begin an iterative process.
Step 3.1:  First, APT-Pipe takes the labeled dataset that was annotated by ChatGPT using the previous Step 2 prompt (without any additional metrics included in the prompt). This offers a baseline to compare results for other prompts that include metrics.
Step 3.2:  Following this, it is necessary to select which metric(s) will be included in the test prompt. To achieve this, we first calculate the metrics for every item in the small manually annotated dataset. For example, if there are four metrics, then each row in the data will be supplemented with four new columns. One option would be to then generate prompts containing every combination of metrics, and test every one to select the best. However, this would be prohibitively slow and expensive. Thus, it is necessary to estimate which metrics are most likely to result in a performance improvement in a more lightweight fashion.
To achieve this, we train an XGBoost classifier (detailed in § 4.5) to estimate the predictive power of each metric. Specifically, we construct a new dataset that contains each data sample, the ground-truth human label, the ChatGPT label (from Step 3.1), and each of the metrics. We then train the classifier to predict if the label generated by ChatGPT will be identical to the original human annotated label, using only ChatGPT’s label, the raw text, and its additional metric (i.e. this is a binary task — either the two labels match, or they do not). Note, we test metrics individually — thus, each metric will have a separately trained classifier. We then calculate the F1 score of each classifier using 10-fold cross-validation, and rank the metrics based on their scores. Our intuition is that metrics that better predict ChatGPT’s ability to generate the correct labels will be more suitable for inclusion in the prompt. This step allows us to reduce the search space and avoid trying every combination of metrics.
Step 3.3:  Next, we select the top-ranked metric and add it to the existing prompt, according to the template below. Using this new test prompt, we ask ChatGPT to annotate the separate validation set and record its proposed labels. This allows us to study if the addition of the metric improves ChatGPT’s performance.
‘‘NLP metrics’’: {
xxxx‘‘Introduction’’: ‘‘Refer to the following NLP metrics of
xxxxthe text to make classification.’’,
xxxx‘‘{metric (e.g., Sentiment)}’’: {
xxxxxxxx{corresponding key-value pairs (detailed in §4.3)}
xxxx},...}
Step 3.4:  If the inclusion of the metric shows a weighted F1-score improvement over the Step 2 baseline, we mark that metric for inclusion in the final prompt.
Iteration:  APT-Pipe then repeats the iteration. We go back to Step 3.1 and re-label the training data using the newly created prompt (which includes the metric marked in Step 3.4 from the previous iteration). For example, if the prior iteration has shown that the inclusion of the toxicity metric improves performance, it is included in the prompt on the next iteration. This re-labeled training data is then used to train again the XGBoost metric classifiers, and select the next metric to add into a new prompt (Step 3.2). Note, only the remaining metrics will be considered, and the ranking of their importance may differ from the previous iteration. Steps 3.3 and 3.4 are then repeated to test if the new metric enhanced ChatGPT’s performance. Note, at this stage, the newly created prompt will include the highest ranked metric, plus all prior metrics that have been shown to improve performance.
Termination Condition:  The loop will continue until all metrics in Step 3.3 have been tested. Thus, the number of iterations is limited by the number of metrics. Importantly, by ranking metrics using XGBoost, we avoid having to experiment with every permutation. Although this cannot be guaranteed to bring optimum performance for all datasets, our experiments have shown that this is an effective estimate. Once all metrics have been tested, APT-Pipe will return a final prompt containing all the chosen metrics as JSON attributes. This becomes the fine-tuned prompt that the prompt engineer will use.
4. Experimental Settings
In this section, we illustrate the selection of social computing datasets for experiment, evaluation metrics, and our pipeline implementation.
4.1. Task Domains and Datasets
Our experiments cover twelve public datasets drawn from three significant domains as highlighted by reviews on text classification (Minaee et al., 2021; Küçük and Can, 2020; Crothers et al., 2023):
  • News classification: Ag’s News (Zhang et al., 2015), SemEval-19 (Kiesel et al., 2019), Sarcasm news headline dataset (Misra and Arora, 2023), and Clickbait news headline dataset (Chakraborty et al., 2016).
  • Stance detection: SemEval-16 (Mohammad et al., 2016), P-Stance (Li et al., 2021), Vax-Stance (Poddar et al., 2022), and RU-Stance (Zhu et al., 2023).
  • AI-writing detection: Tweepfake (Fagni et al., 2021), FakeHate (De Gibert et al., 2018; Hartvigsen et al., 2022), GPT-Article (Maktab Dar Oghaz et al., 2023), and GPT-Wiki (Holley, 2023).
We summarize these datasets’ statistics and descriptions in Table 7. For any dataset with more than 3,000 items, we subset it to 3,000 items using stratified sampling by its classification labels. We do this due to the rate and cost limitations of the ChatGPT API.
4.2. Evaluation Metrics
We assess the tuned prompts’ performance by our pipeline based on the following evaluation metrics:
  • Classifier performance: We view ChatGPT as a text classification engine and use standard classifier metrics: weighted F1-score, precision, and recall.
  • Responses’ parsability: The percentage of parsable responses based on the desired format. We assume a response is parsable only if it reports the label in the desired format. We use regular expressions to match a string in JSON syntax from ChatGPT’s response and then parse it into a JSON object.
  • Time cost: The time it takes for ChatGPT to return an annotation response. We normalize the request time by comparing it to the request time of a null prompt sent before. Here, the normalization is done to remove the influence of the network transmission. Afterwords, we divide the normalized request time by the number of words in the response. This metric can be interpreted as the time required for ChatGPT to generate each token.
4.3. NLP Metrics for Prompt Augmentation
In Step 3, we automatically select metadata metrics to include in the prompt. For the purpose of our evaluation, we experiment with four examples of NLP metrics. Table 4.3 summarizes how we include the metrics in each prompt. The metrics are:
  • Sentiment: This metric indicates the text’s sentiment polarity. We use a pre-trained sentiment analysis model called XLM-T to measure the sentimental polarity of the text using a three-dimensional vector represented as positive, negative, and neutral (Barbieri et al., 2021). For prompt augmentation, we input sentimental polarity (ranging from 0 to 1) based on above three dimensions.
  • Emotion: This metric indicates the text’s emotional leaning. We utilize a well-known model called Emotion English DistilRoBERTa-base, which uses a seven-dimensional vector, covering Ekman’s six basic emotions with a neutral class (Ekman et al., 1999; Hartmann, 2022). For prompt augmentation, we input emotional leaning (ranging from 0 to 1) for the above seven listed emotions.
  • Toxicity: This metric indicates the percentage of toxic content contained in the text. We use the Google Perspective API to measure the existence of toxic content. We construct a six-dimensional vector covering the six Perspective attributes (Jigsaw, 2017) used to measure toxicity (e.g., TOXICITY, IDENTITY_ATTACK). For prompt augmentation, we input the API’s scores (ranging from 0 to 1) for these six attributes.
  • Topic: This metric indicates the main topics of the text. Given a text, we train a BERTopic model using the dataset it belongs to and then infer its topic embeddings according to the trained model (Grootendorst, 2022). For prompt augmentation, we extract the most representative topic as the one with the highest possibility in its BERTopic embedding. We then input into the prompt the list of important words of this topic reported by the model. Note, alternative configurations could include a larger number of topics.
Prompt Augmentation
Sentiment ‘‘Introduction’’: ‘‘Scores of sentiment leaning of text (ranging from 0 to 1).’’,
“Scores”: {“Positive”: 0.xx, “Neutral”: 0.xx, “Negative”: 0.xx}
Emotion
‘‘Introduction’’: ‘‘Scores of emotion leaning of text (ranging from 0 to 1).’’,
“Scores”: {“Anger”: 0.xx, “Disgust”: 0.xx, “Fear”: 0.xx, “Joy”: 0.xx, “Neutral”: 0.xx, “Sadness”: 0.xx, “Surprise”: 0.xx}
Toxicity
‘‘Introduction’’: ‘‘Scores of toxcity degree of text (ranging from 0 to 1).’’,
“Scores”: {“Overall Toxicity”: 0.xx, “Severe Toxicity”: 0.xx, “Identity Attack”: 0.xx, “Insult”: 0.xx, “Profanity”: 0.xx, “Threat”: 0.xx}
Topic
‘‘Introduction’’: ‘‘Representative words to describe the major topic of the text.’’,
“Words”: [keywords to describe text’s major topic by BERTopic]
Table 1. A summary of key-value pairs augmented in the prompt text for prompt-tuning on NLP metrics. The blue text denotes the probability or text output by the corresponding model.
4.4. Baselines
We select several basline prompts generated by existing tuned templates to compare with the APT-Pipe generated prompts.
Cloze prompt:  This template designs prompts as a cloze-style question. It asks ChatGPT to replace the blank or masked part of the text. Prior studies have utilized cloze prompts to improve ChatGPT’s performance on text classification (Xiang et al., 2022; Han et al., 2022). We generalize a cloze prompt template as follows:
Fill [Label] for {task domain} task with a label in [Label 1, Label 2, ...].
The text ‘‘{text to classify}’’ is classified as [Label].
Desired format:
Label: <label_for_classification>
Dictionary prompt:  This template is a generalized pattern representing prompts formed as a dictionary. These are commonly applied in text classification studies on ChatGPT (Zhu et al., 2023; Bang et al., 2023; Ding et al., 2022). We generalize a dictionary prompt as follows:
Classify the following text by given labels for specified task.
Text: ‘‘{text to classify}’’.
Task: {task domain}.
Labels: [Label 1, Label 2, ...].
Desired format:
Label: <label_for_classification>
JSON prompt:  This template represents the prompts encoded in JSON format in Step 1, without any further prompt-tuning (§ 3.1).
Parsing responses:  When each baseline prompt generates a response, it is necessary to parse it and extract the returned annotation label. For responses generated by the dictionary and cloze prompts, we index ChatGPT’s classification label baseline as the first word of the substring after “label:”. For responses generated by JSON prompts, we parse them into a JSON object and extract ChatGPT’s classification label using the key “label”
4.5. Experimental Implementation
ChatGPT setting:  We have built the full APT-Pipe tool, and integrated it with gpt-3.5-turbo. We request access ChatGPT through OpenAI’s API with parameter temperature set to 00 to make the response focused and offer deterministic. [^1] https://platform.openai.com/docs/api-reference/chat APT-Pipe will be made publicly available on github. [^2] Redacted due to anonymous submission requirements.
Dataset Split:  The experiment use a data split strategy where we divide each dataset into training, validation, and test sets, allocated at a ratio of 60%, 20%, and 20%, respectively. Initially, in Step 1, we merge the training and validation sets to create a single input dataset for APT-Pipe (covering 80%). Subsequently, in Step 3, APT-Pipe separates them again and uses the test and validation sets separately for the prompt-tuning with NLP metrics. Thus, we work on the assumption that users of APT-Pipe will be able to annotate a relatively small number of annotations to drive APT-Pipe’s tuning. Note that the remaining 20% test set is not used within APT-Pipe and instead is used only for evaluation purposes in Section 5.
Classifier for metric ranking:  In Step 3, we train a classifier for each NLP metric for ranking and selecting the metrics for prompt augmentation. For this, we use Extreme Gradient Boosting (XGBoost) (Chen and Guestrin, 2016). We implement it using the xgboost python package, with parameter setting (objective=“binary:logistic”, seed=42).
5. Results
We next evaluate APT-Pipe’s ability to improve prompt performance. We further perform an ablation study to show the effectiveness of its stepwise prompt-tuning methods.
5.1. Overall Annotation Performance
Shot NLP Metrics (in selection order) F1-score Precision Recall
Cloze Dictionary JSON APT-Pipe Cloze Dictionary JSON APT-Pipe Cloze Dictionary JSON APT-Pipe
[…]
Sarcasm Zero Emotion. 61.24% 62.46% 52.34% 55.31% 67.95% 68.35% 65.40% 66.88% 63.57% 64.47% 58.01% 59.87%
[…]
P-Stance Zero - 79.69% 79.10% 77.86% 77.86% 80.61% 80.74% 81.02% 81.02% 79.86% 79.28% 78.41% 78.41%
[…]
GPT-Article Few Sentiment, Emotion, Toxicity 60.40% 46.87% 54.05% 40.96% 61.21% 50.93% 58.72% 48.92% 60.78% 52.63% 56.78% 49.75%
[…]
We first compare APT-Pipe’s performance against the three baselines. Table 2 presents ChatGPT’s weighted F1-score, precision, and recall on the test sets of all twelve datasets. Overall, APT-Pipe achieves the highest weighted F1-score on nine datasets. Compared with the baseline with the second-best F1-score, APT-Pipe improves the F1-score by 7.01% ($SD=9.74\%$) on average. This evidences that APT-Pipe is effective in improving ChatGPT’s overall annotation performance. On average, APT-Pipe also improves ChatGPT’s weighted precision by 1.18% ($SD=1.36\%$) for eight datasets, and by 4.88% ($SD=7.70\%$) for the weighted recall rate for nine datasets. This indicates that APT-Pipe also enables ChatGPT to respond with accurate annotations with fewer false positive answers.
Recall that each task (news, stance and AI-writing classification) has four evaluation datasets. APT-Pipe performs the best on 3/4 of their datasets for each task. Thus, we confirm that APT-Pipe can benefit ChatGPT on annotating social computing data for news classification, stance detection, and AI-writing detection. However, the improvements in ChatGPT’s performance vary across these three tasks. For news classification and stance detection, APT-Pipe offers an average improvement of 2.20% ($SD=0.80\%$) and 1.22% ($SD=0.48\%$) F1-score, respectively. APT-Pipe’s improvements are far greater for AI-writing detection though, with an average improvement of 17.61% F1-score ($SD=11.19\%$). Such results suggest that, although APT-Pipe can enhance ChatGPT’s annotation on social computing data, its effectiveness depends on the task domain. In our case, APT-Pipe has a much larger advantage in AI-writing detection than news classification and stance detection.
5.2. Effectiveness of ChatGPT’s Responses
Responses’ parsability Time cost
Cloze Dictionary JSON APT-Pipe Cloze Dictionary JSON APT-Pipe
Ag’s News 98.98% 57.69% 99.83% 99.83% 0.439 0.546 0.197 0.250
SemEval-19 100.00% 93.60% 100.00% 100.00% 0.246 0.272 0.154 0.168
Sarcasm 100.00% 98.29% 100.00% 100.00% 0.264 0.304 0.169 0.172
Clickbait 100.00% 66.89% 100.00% 99.83% 0.231 0.279 0.160 0.199
SemEval-16 89.30% 65.01% 93.01% 94.68% 0.287 0.253 0.156 0.162
P-Stance 98.63% 56.53% 96.42% 96.42% 0.247 0.175 0.168 0.168
Vax-Stance 100.00% 100.00% 100.00% 96.52% 0.360 0.492 0.196 0.146
RU-Stance 100.00% 97.57% 98.95% 97.55% 0.280 0.374 0.172 0.176
Tweepfake 100.00% 81.99% 100.00% 96.76% 0.468 0.112 0.199 0.191
FakeHate 100.00% 81.57% 100.00% 91.80% 0.546 0.123 0.203 0.227
GPT-Article 100.00% 68.21% 100.00% 98.53% 0.478 0.079 0.199 0.263
GPT-Wiki 100.00% 100.00% 100.00% 92.98% 0.782 0.736 0.294 0.629
Average 98.91% 80.61% 99.02% 97.08% 0.386 0.312 0.188 0.229
(Std) (3.06%) (17.09%) (2.16%) (2.79%) (0.165) (0.197) (0.038) (0.131)
Table 3. Comparison between the effectiveness on ChatGPT’s responses by baseline prompts and APT-Pipe prompts.
As implied in relevant studies, two challenges for applying prompt-tuning for data annotation tasks are that (i) ChatGPT may generate responses in arbitrary formats (Amin et al., 2023; Guo et al., 2023); and (ii) tuning methods could increase the time cost for generating annotations (Shi and Lipani, 2023). We next assess whether APT-Pipe also faces these two difficulties.
Parsability:  Table 3 shows responses’ parsability for ChatGPT on all 12 dataset. On average, 97.08% ($SD=2.79\%$) of responses generated by APT-Pipe are parsable. When compared with baselines, APT-Pipe improves upon dictionary prompts, with an improvement of 16.47% parsability on average. However, the parsability is slightly reduced compared to the cloze or plain JSON prompts. Applying APT-Pipe decreases the responses’ parsability by 1.83% when compared to cloze prompts, and 1.94% when compared to JSON prompts on average. These results show that applying APT-Pipe would pay a small price of responses’ parsability. We conjecture this is because adding key-value pairs for prompt-tuning also increases the entropy of the prompt text, which can lead LLMs to respond to answers with a higher uncertainty (Yu et al., 2023b; Borji, 2023).
Time Cost:  Table 3 also presents the time taken for ChatGPT to return an annotation. We see that APT-Pipe improves upon both cloze and dictionary prompts. Applying APT-Pipe decreases the time cost by 0.157 seconds per token when compared to cloze prompts, and 0.082 seconds per token when compared to plain JSON prompts. However, compared with JSON prompts, APT-Pipe introduces a small additional time cost (0.041 seconds per token). These results show that the tuned prompts take slightly longer to execute compared to the un-tuned JSON. This is because the tuned prompts contain additional information, which is used to achieve performance improvement. Nonetheless, APT-Pipe is still able to boost ChatGPT’s speed for data annotation compared to the dictionary and cloze prompts utilized in other literature.
5.3. Ablation Study
F1-score Precision Recall
APT-Pipe w/o Step 2 w/o Step 3 APT-Pipe w/o Step 2 w/o Step 3 APT-Pipe w/o Step 2 w/o Step 3
Ag’s News 88.32% 83.43% 88.74% 88.72% 85.70% 88.84% 88.31% 83.87% 88.76%
SemEval-19 80.29% - 77.47% 81.08% - 78.00% 80.00% - 77.27%
Sarcasm 55.31% - 51.54% 66.88% - 69.28% 59.87% - 58.38%
Clickbait 94.56% - 93.90% 94.64% - 94.25% 94.57% - 93.92%
SemEval-16 50.93% 34.90% 45.10% 57.95% 52.37% 53.43% 47.28% 31.33% 41.53%
P-Stance 74.82% - - 77.27% - - 75.33% - -
Vax-Stance 66.25% - 63.11% 69.97% - 65.50% 64.91% - 62.81%
RU-Stance 74.89% - 73.17% 81.58% - 80.98% 74.55% - 72.79%
Tweepfake 75.27% 62.40% - 76.20% 70.81% - 75.46% 64.86% -
FakeHate 49.51% 24.53% - 74.56% 67.35% - 49.18% 31.54% -
GPT-Article 40.96% 51.79% 42.75% 48.92% 58.70% 43.15% 49.75% 56.00% 43.13%
GPT-Wiki 65.11% 39.54% - 67.17% 67.13% - 64.15% 48.33% -
Table 4. Ablation study on prompt-tuning. The “w/o Step 2/3” represents the variant of APT-Pipe without prompt-tuning in Step 2/3. A “-” denotes that the ablation comparison is not available as APT-Pipe prompts on the corresponding dataset don’t apply few-shot learning in Step 2 or no NLP metrics are selected in Step 3. The bold percentage denotes the highest value of corresponding metrics.
We next conduct an ablation study to measure the improvements attained by each step described in Section 3. Note, Table 2 already shows the outputs of Step 1 in the JSON column. Thus, we set up two additional variant pipelines, and compare them against APT-Pipe:
  • Without Step 2: In this variant pipeline, we discard Step 2 from APT-Pipe. We directly apply the NLP feature selection module (Step 3) to the initialized JSON prompt (from Step 1), and then add the selected NLP features to the encoded prompt.
  • Without Step 3: In this variant pipeline, we remove step 3 from APT-Pipe. It only uses the few shots or the zero shots prompting mechanism from Step 2.
Table 4 compares ChatGPT’s performance on the test sets using the original pipeline, and the two variants. APT-Pipe outperforms the variants in nine out of twelve datasets in terms of F1 scores. We observe that APT-Pipe outperforms the Without Step 2 method by 6.12% on F1 score, 0.96% on precision, and 4.85% on recall. For the Without Step 3 method, APT-Pipeline outperforms it by 1.31% on F1 score, 1.36% on precision, and 1.72% on recall. This demonstrate the effectiveness of our fully integrated prompting steps in improving the performance of ChatGPT for text annotation tasks.
6. Extending APT-Pipe Prompts
Table 5. A summary of key-value pairs augmented in prompt text in APT-Pipe prompts for applying Chain-of-Thought (CoT) and Tree-of-Thought (ToT).
Zero-shot Few-shot
CoT
‘‘Thought’’: ‘‘Let’s think step by step.’’ ‘‘Thought’’: ‘‘Let’s think step by step.’’,
“Examples for thought”: [{“Text”: “{exemplar 1}”, “Label”: “{exemplar 1’s ground-truth label}”}, “Explanation”: “{CoT explanation for exemplar 1’s ground-truth label}”},…]
ToT
‘‘Thought’’: ‘‘Imagine three different experts are answering this question. All experts will write down 1 step of their thinking, then share it with the group. Then all experts will go on to the next step, etc. If any expert realises they’re wrong at any point then they leave. Finally, all experts vote and elect the majority label as the final result.’’ ‘‘Thought’’: ‘‘Imagine three different experts are answering this question. All experts will write down 1 step of their thinking, then share it with the group. Then all experts will go on to the next step, etc. If any expert realises they’re wrong at any point then they leave. Finally, all experts vote and elect the majority label as the final result.’’,
“Examples for thought”: [{“Text”: “{exemplar 1}”, “Label”: “{exemplar 1’s ground-truth label}”, “Explanation”: “{ToT explanation for exemplar 1’s ground-truth label}”}, …]
F1-score Precision Recall
w/i CoT w/i ToT APT-Pipe w/i CoT w/i ToT APT-Pipe w/i CoT w/i ToT APT-Pipe
Ag’s News 84.26% 85.60% 88.32% 86.75% 86.90% 88.72% 84.59% 85.79% 88.31%
[…]
Clickbait 95.43% $\uparrow$ 92.71% 95.35% 95.52% $\uparrow$ 93.42% 95.39% 95.43% $\uparrow$ 92.73% 95.34%
SemEval-16 45.08% 45.42% 50.93% 53.21% 53.16% 57.95% 42.14% 42.64% 47.28%
P-Stance 76.81% 80.52% $\uparrow$ 77.86% 79.95% 82.22% $\uparrow$ 81.02% 77.43% 80.78% $\uparrow$ 78.41%
Vax-Stance 68.04% $\uparrow$ 59.50% 65.58% 71.54% $\uparrow$ 60.25% 69.31% 67.00% $\uparrow$ 58.96% 64.26%
RU-Stance 74.27% 76.25% $\uparrow$ 74.89% 81.97% $\uparrow$ 81.12% 81.58% 73.87% 75.66% $\uparrow$ 74.55%
Tweepfake 72.52% 71.40% 75.27% 73.32% 74.68% 76.20% 72.68% 72.05% 75.46%
FakeHate 49.18% 20.99% 49.51% 71.61% 71.43% 74.56% 50.41% $\uparrow$ 30.29% 49.18%
GPT-Article 55.04% $\uparrow$ 49.86% $\uparrow$ 40.96% 55.43% $\uparrow$ 50.06% $\uparrow$ 48.92% 54.97% $\uparrow$ 50.31% $\uparrow$ 49.75%
GPT-Wiki 67.23% $\uparrow$ 70.58% $\uparrow$ 65.11% 69.41% $\uparrow$ 76.31% $\uparrow$ 67.17% 66.67% $\uparrow$ 69.81% $\uparrow$ 64.15%
Table 6. Comparison between ChatGPT’s performance within APT-Pipe prompts extended with CoT and ToT. The “w/i CoT (ToT)” represents the group of prompts by APT-Pipe extended with CoT (ToT). A $\uparrow$” beside a percentage means CoT/ToT improves APT-Pipe’s performance. The bold percentage denotes the highest value of corresponding metrics.
APT-Pipe is built in a modular fashion and is extensible by design. To demonstrate this, we briefly conduct a case-study by extending APT-Pipe to support two additional state-of-art tuning mechanisms:
Implementation:  For this experiment, we “plug-in” these two modules to create a fourth stage in the pipeline. Here, we embed the two proposed augmentations into the prompts output by APT-Pipe in Step 3 (Kojima et al., 2022; Hulbert, 2023). Both augmentations are formed by adding instructive sentences that can be directly injected into prompts to guide ChatGPT’s reasoning process. Table 6 lists the specific text for the two prompt additions. For CoT, its augmentation involves the simple addition of the instruction “Let’s think step by step.” (Kojima et al., 2022). For ToT, its augmentation involves more comprehensive guidance as “Imagine three different experts are answering this question. All experts will write down 1 step of their thinking, then share it with the group. Then all experts will go on to the next step, etc. If any expert realises they’re wrong at any point then they leave. Finally, all experts vote and elect the majority label as the final result.” (Hulbert, 2023).
Note that prompts tuned by APT-Pipe may employ either zero-shot or few-shot configurations. We refer to suggestions by relevant studies to set up the two types of automatic realizations of CoT/ToT under zero-shot and few-shot settings, respectively (Shum et al., 2023; Long, 2023). For APT-Prompts with a zero-shot configuration, applying CoT and ToT simply adds their instructive sentences into the prompts (see Table 6). For APT-Pipe prompts with the few-shot configuration, for each prompt and its exemplars selected in Step 2, we ask ChatGPT to explain the exemplars’ ground-truth labels through a prompt enhanced by CoT/ToT. Afterwards, we append these explanations to the exemplars in prompts correspondingly. The CoT/ToT prompt enhancement is as follows:
Results:  Table 6 presents the performance on the test sets using the baseline APT-Pipe prompts (from Section 3) against their extended versions with CoT and ToT. CoT improves APT-Pipe’s F1-scores on five datasets by 4.01% ($SD=5.71\%$) on average, and ToT improves APT-Pipe’s F1-scores on six datasets by 3.49% ($SD=3.17\%$) on average. These results show that extending APT-Pipe with CoT and ToT has the potential to improve performance. However, we also identify some trade-offs for such extensions. For example, ToT decreases ChatGPT’s precision on eight datasets by 3.84% ($SD=3.16\%$) on average. This implies that extending APT-Pipe with ToT can make ChatGPT less likely to respond with precise annotations and introduce more false positive answers instead. Importantly, APT-Pipe can evaluate the efficacy for any given dataset and only include the prompt additions that result in performance improvements. We, therefore, show this to highlight APT-Pipe’s modular flexibility.
Table 5. A summary of key-value pairs augmented in prompt text in APT-Pipe prompts for applying Chain-of-Thought (CoT) and Tree-of-Thought (ToT).